# Astute RAG: Overcoming Imperfect Retrieval Augmentation and Knowledge Conflicts for Large Language Models

### v1

@@ 3.4 Iterative Source-aware Knowledge Consolidation
We additionally ensure source-awareness by providing the source of each passage to LLMs when consolidating knowledge. The source information (internal or external, such as a website) is helpful in assessing the reliability of passages. Here, we provide the passage source as $S_{0}=[\mathbbm{1}_{\{d\in E\}}\text{for}\ d\ \text{in}\ D_{0}].$
To consolidate knowledge, we prompt the LLM (with $p_{con}$ in Appendix A) to identify consistent information across passages, detect conflicting information between each group of consistent passages, and filter out irrelevant information. This step would regroup the unreliable knowledge in input passages into fewer refined passages. The regrouped passages will also attribute their source to the corresponding one or more input passages
@@
We find that this is especially helpful in comparing the reliability of conflicting knowledge and addressing knowledge conflicts. Moreover, this knowledge consolidation process can run iteratively for $t$ times to improve the context to be more and more useful. Users can assign a larger number of iterations when the context is lengthy.
@@ 3.5 Answer Finalization
In the last step, we prompt the LLM (with $p_{ans}$ in Appendix A) to generate one answer based on each group of passages ($\langle D_{t},S_{t}\rangle$), and then compare their reliability and select the most reliable one as the final answer. This comparison allows the LLM to comprehensively consider knowledge source, cross-source confirmation, frequency, and information thoroughness when making the final decision. Notably, this step can be merged into the last knowledge consolidation step to reduce the inference complexity (the amount of prediction API calls) using a combined prompt:
@@
When $t=1$, the initial passages will be fed into the model directly for knowledge consolidation and subsequent answering: $a=\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle).$
@@ 4 Experiments
We evaluate the effectiveness of Astute RAG on overcoming imperfect retrieval augmentation and addressing knowledge conflicts. In this section, we first introduce the experiment setting in detail (Section 4.1). Then, we compare the performance of Astute RAG with various baselines on diverse datasets (Section 4.2). Finally, we provide in-depth analyses (Section 4.3).
Method #API Calls NQ TriviaQA BioASQ PopQA Overall
Claude 3.5 Sonnet (20240620)
No RAG 1 47.12 81.98 50.35 29.78 54.51
RAG 1 44.41 76.68 58.04 35.96 55.47
USC (Chen et al., 2024b) 4 48.14 80.21 61.54 37.64 58.73
GenRead (Yu et al., 2023a) 2 42.03 74.20 56.99 34.27 53.55
RobustRAG (Xiang et al., 2024) 11 47.80 78.09 56.29 37.08 56.53
InstructRAG (Wei et al., 2024) 1 47.12 83.04 58.04 41.01 58.83
Self-Route (Xu et al., 2024a) 1-2 47.46 78.80 59.09 41.01 58.06
Astute RAG (t=1) 2 52.20 84.10 60.14 44.38 61.71
Astute RAG (t=2) 3 53.22 84.45 61.89 44.94 62.67
Astute RAG (t=3) 4 53.56 84.45 62.24 44.94 62.86
Table 1: Main results on Claude under zero-shot setting, showing the accuracy of different benchmark methods vs. Astute RAG, along with their prediction complexity, in number of prediction API calls. Best scores are in bold.
@@ 4.1 Experimental Settings
Datasets and metrics. We conduct experiments on the data collected in Section 2 consisting of data from NQ, TriviaQA, BioASQ, and PopQA. For each instance from these datasets, we provide 10 passages collected under a realistic retrieval setting: for each question in our benchmark, we query Google Search to retrieve the top 30 results and select the first 10 accessible websites. From each retrieved website, we extract the paragraph corresponding to the snippet provided in Google Search results as the retrieved passage.. Most of the retrieval results contains natural noise with irrelevant or misleading information. We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems. Notably, we do not select questions or annotate answers based on the retrieval results. This setting allows us to analyze the severity of imperfect retrieval in real-world RAG. It distinguishes our benchmark from previous ones that employ synthetic retrieval corruptions or that unintentionally reduce the frequency of imperfect retrieval with biased construction protocols (Chen et al., 2024a; Yang et al., 2024). We also evaluate our method on RGB (Chen et al., 2024a), a RAG diagnostic benchmark evaluating several crucial RAG abilities. Specifically, we choose the English subset of RGB focusing on noise robustness. The benchmark have positive and negative passage sets for each question. We select five negative documents per question as the context to form a worst-case scenario. All the data in these datasets are short-form QA. Following previous work (Xiang et al., 2024; Wei et al., 2024; Mallen et al., 2023), a model response is considered correct if it contains the ground-truth answer. To enhance evaluation reliability, we prompt LLMs to enclose the exact answer within special tokens, extracting them as the final responses.
@@
General Settings of LLMs and RAG. We conduct experiments on two advanced LLMs, including Gemini 1.5 Pro [^5] https://deepmind.google/technologies/gemini/pro/ (gemini-1.5-pro-002) and Claude 3.5 Sonnet [^6] https://www.anthropic.com/news/claude-3-5-sonnet (claude-3-5-sonnet@20240620). The generation temperature is set to 0 and the maximum output tokens is set to 1,024, if not specified otherwise. By default, the passages are presented in the prompt by reversed order. All experiments are under the zero-shot setting for controlled evaluation, where no demonstrations for QA or method-specific steps are provided.
@@
Baselines. We compare Astute RAG with various RAG methods designed for enhanced robustness and representative inference strategies designed to improve response trustworthiness. USC (Chen et al., 2024b) is the universal self-consistency method that samples multiple LLM responses given the same context and aggregates the answers. It provides a reference of naive improvements using additional API calls. The temperature for sampling responses in this baseline is set to 0.7. Genread (Yu et al., 2023a) augments retrieved passages with LLM-generated passages. It provide a reference of presenting passages from both internal and external knowledge in the prompt without effectively combining them. RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness. We use the keyword aggregation variant as it is shown to be the best-performing variant on advanced LLMs. InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages. For a fair comparison, we use the instructions without training or in-context learning. Self-Route (Xu et al., 2024a) adaptively switches between LLMs with and without RAG. [^7] The original Self-Route switches between RAG and long-context LLMs, while our implementation switches between LLMs with and without RAG to better align with the problem formulation in this paper. This baseline provides a reference of switching between LLMs’ internal and external knowledge.
Implementation Details of Astute RAG. The prompt templates for Astute RAG can be found in Appendix A. By default, we use 2 API calls per query, setting $t=1$ to merge the prompt for knowledge consolidation and answer finalization. For adaptive generation of internal knowledge, we prompt the LLM to generate no more than one passage.
Method #API Calls NQ TriviaQA BioASQ PopQA Overall
Gemini 1.5 Pro (002)
No RAG 1 44.75 80.21 45.80 25.28 51.34
RAG 1 42.71 75.97 55.24 33.71 53.65
USC (Chen et al., 2024b) 4 46.44 76.68 58.39 37.64 56.43
GenRead (Yu et al., 2023a) 2 45.08 77.39 54.90 34.27 54.70
RobustRAG (Xiang et al., 2024) [^8] We observe a high refusal rate in responses of RobustRAG. 11 34.24 67.49 44.06 32.02 45.59
InstructRAG (Wei et al., 2024) 1 46.78 80.57 54.90 34.83 56.14
Self-Route (Xu et al., 2024a) 1-2 47.46 79.86 58.04 38.20 57.58
Astute RAG (t=1) 2 50.17 81.63 58.04 40.45 59.21
Astute RAG (t=2) 3 51.53 81.27 58.74 40.45 59.69
Astute RAG (t=3) 4 48.47 80.21 60.14 42.13 59.21
Table 2: Main results on Gemini under zero-shot setting, showing the accuracy of different benchmark methods vs. Astute RAG, along with their prediction complexity, in number of prediction API calls. Best scores are in bold.
@@ 4.2 Main Results
Table 1 and Table 2 presents the results on data with realistic retrieval augmentation for each dataset. By comparing RAG and No RAG, we find that retrieved passages might not always bring benefits to downstream performance – on NQ and TriviaQA, RAG performance lags behind No RAG. We attribute this to that the questions being covered by the LLM’s internal knowledge and the noise in retrieval results misleading the LLM. In contrast, on BioASQ and PopQA, which focus on domain-specific and long-tail questions, RAG significantly improves LLM performance. However, due to imperfect retrieval augmentation, the absolute performance still remains to be unsatisfactory. Among all baselines, no single method consistently outperforms others across all datasets. This observation highlights that these baselines are tailored to distinct settings and may not be universally applicable. For instance, InstructRAG is more effective on TriviaQA, achieving the best performance among all baselines with both Claude and Gemini. In contrast, Self-Route performs better than InstructRAG on both NQ and BioASQ. Moreover, RobustRAG achieves very different performance when applied to Gemini and Claude. Through in-depth analysis, we find that RobustRAG with Gemini exhibits a high refusal rate (refuse to answer) in responses. We attribute this instability to the varying method designs of the baselines, which are tailored for different scenarios, resulting in inconsistent improvement across datasets. Overall, InstructRAG and Self-Route demonstrates the best performance among all baselines when applied to Claude and Gemini respectively. We also note that increasing the number of API calls does not necessarily correlate with improved performance.
Astute RAG consistently outperforms baselines across all datasets of different properties. The overall improvement compared with the best baseline is relatively 6.85% on Claude and 4.13% on Gemini, and the improvements in domain-specific questions are much higher. These results highlight the effectiveness of Astute RAG in overcoming imperfect retrieval augmentation. On Claude, adding more iteration of knowledge consolidation leads to consist improvement. The improvement margin becomes lower when $t$ becomes larger. This is because after each iteration, the remaining improvement space for knowledge consolidation becomes smaller. On Gemini, increasing $t$ primarily benefits BioASQ and PopQA. These two datasets rely more heavily on external knowledge, and iterative knowledge consolidation helps mitigate noise within this external information. Performance on NQ and TriviaQA does not improve further when $t$ reaches 3. We attribute this to the less critical role of external knowledge in these datasets. For setting consistency and efficiency, we set the parameter $\hat{m}$ to a smaller value, limiting the influence of internal knowledge.
@@ 4.3 Analyses
Performance by retrieval precision. We compare the performance of Astute RAG and baselines across different subsets partitioned by their retrieval precision, on our collected data with Claude as the LLM. As shown in Figure 4, Astute RAG achieves consistently better performance than all baselines across different retrieval precision, indicating its effectiveness in improving RAG trustworthiness in broad scenarios. Notably, Astute RAG does not sacrifice performance gain under high retrieval quality in exchange for improvement under low retrieval quality. When the retrieval quality is extremely low (close to zero retrieval precision), all other RAG variants underperforms the ’No RAG’ baseline, except for the proposed Astute RAG. This observation aligns with the worst-case results on RGB. It demonstrates the difficulty in overcoming imperfect retrieval augmentation, and verify the effectiveness of Astute RAG in doing so.
(a) Performance by Retrieval Precision
(b) Performance by Knowledge Conflicts
(c) Worst-case Performance on RGB
Figure 4: Analysis results. (a) Performance of Claude across different buckets ranked by retrieval precision. (b) Performance of Claude on conflicting and consistent instances between No RAG and RAG. (c) Worst-case performance of Claude on RGB, where all retrieved passages are negative. Astute RAG reaches a performance close to No RAG, while other RAG systems are far behind.
@@
Effectiveness in addressing knowledge conflicts. We split our collected data in to three subset according to the answers from Claude, with and without RAG. The answers from two inference methods can be both correct, both incorrect, or conflicting with one being correct. These three subsets represents the three situations between internal and external knowledge. The results are shown in Figure 4. On the conflicting subset, Astute RAG successfully chooses the correct answer in approximately 80% of cases, being the most effective method in addressing knowledge conflicts. Notably, Astute RAG even brings performance improvement on the subset where neither internal nor external knowledge alone leads to the correct answer. This indicates that Astute RAG can effectively combine partially-correct information from LLM-internal and external knowledge, to achieve the correct answer through collective information across them.
Worst-case performance on RGB.. Figure 4 presents the results under the worst-case setting on RGB where all retrieved documents are negative. It demonstrates the noise robustness of Astute RAG and baseline RAG methods. The performance gap between RAG and No RAG exceeds 50 points, highlighting the detrimental impact of imperfect retrieval results and emphasizing the importance of providing robust safeguards against worst-case scenarios. While the baseline RAG methods outperform the original RAG, they still obviously fall behind No RAG. Astute RAG is the only RAG method that reaches a performance close to No RAG under the worst-case scenario, further supporting its effectiveness in addressing imperfect retrieval augmentation.
@@
Qualitative study. In Figure 5, we present two representative examples showing the intermediate outputs of Astute RAG. In the first example, LLM without RAG generates a wrong answer, while RAG returns a correct answer. Astute RAG successfully identified the incorrect information in its generated passage and an external passage, avoiding confirmation bias Tan et al. (2024). In the second example, LLM alone is correct, while RAG is incorrect due to the noisy retrieval results. Astute RAG detected the correct answer from noisy retrieved information by checking with its internal knowledge.
Figure 5: Qualitative examples. Top: Astute RAG identified the error in internal knowledge (i.e., generated passage) by confirming with external sources. Bottom: Astute RAG detected the correct answer from noisy retrieved information by checking with its internal knowledge. Standard RAG does not provide an answer because the retrieved passages are too noisy.
5 Related Work
Retrieval augmented generation (RAG) seeks to address the inherent knowledge limitation of LLMs with passages retrieved from external sources of information such as private corpora or public knowledge bases (Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022). Given the widespread adoption of RAG in various real-world applications, including risk-sensitive domains, the negative impact of noisy information within retrieved passages has garnered increasing attention (Cuconasu et al., 2024). Recent work has sought to enhance the robustness of RAG systems against noise from various perspectives, including training LLMs with noisy context (Yu et al., 2023b; Yoran et al., 2024; Pan et al., 2024; Fang et al., 2024), training small models to filter out irrelevant passages (Wang et al., 2023c; Xu et al., 2023), passage reranking (Yu et al., 2024; Glass et al., 2022), dynamic and iterative retrieval (Jiang et al., 2023; Asai et al., 2023; Yan et al., 2024), query rewriting (Ma et al., 2023), and speculative drafting (Wang et al., 2024). These studies focus on distinct modules or stages of RAG systems and are orthogonal to our work.
Our work focuses on enhancing RAG robustness at the post-retrieval stage, after retrieved passages have been provided. On this topic, RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness. InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages. MADRA (Wang et al., 2023b) applies multi-agent debate to select helpful evidence. However, these works do not explicitly incorporate internal knowledge to recover from RAG failures and may therefore collapse when the majority of retrieved passages are negative. In terms of emphasizing internal knowledge of LLMs in RAG, recent work has explored using LLM-generated passage as context (Yu et al., 2023a; Zhang et al., 2023), adaptively switching between LLMs with and without RAG (Xu et al., 2024a; Mallen et al., 2023; Jeong et al., 2024), and combining answers from internal and external knowledge through contrastive decoding (Zhao et al., 2024; Jin et al., 2024). We focus on a black-box setting where no further training is required, directly addressing knowledge conflicts to combine the helpful information from both sides and achieve more reliable answers.

### v2

@@ 4.3 Iterative Source-aware Knowledge Consolidation
We additionally ensure source-awareness by providing the source of each passage to LLMs when consolidating knowledge. The source information (internal or external, such as a website) is helpful in assessing the reliability of passages. Here, we provide the passage source as $S_{0}=[\mathbbm{1}_{\{d\in E\}}\text{for}\ d\ \text{in}\ D_{0}].$ To consolidate knowledge, we prompt the LLM (with $p_{con}$ in Appx. A) to identify consistent information across passages, detect conflicting information between each group of consistent passages, and filter out irrelevant information. This step would regroup the unreliable knowledge in input passages into fewer refined passages. The regrouped passages also attribute their source to the corresponding input passages:
@@
We find that this is especially helpful in comparing the reliability of conflicting knowledge and addressing knowledge conflicts. This knowledge consolidation process can run iteratively for $t$ times to improve better utilization of the retrieved context.
@@ 4.4 Answer Finalization
In the last step, we prompt the LLM (with $p_{ans}$ in Appx. A) to generate one answer based on each group of passages ($\langle D_{t},S_{t}\rangle$), and then compare their reliability and select the most reliable one as the final answer. This comparison allows the LLM to comprehensively consider knowledge source, cross-source confirmation, frequency, and information thoroughness when making the final decision. Notably, this step can be merged into the last knowledge consolidation step to reduce the inference complexity (the amount of prediction API calls) using a combined prompt:
@@
When $t=1$, the initial passages will be input to the model directly for knowledge consolidation and subsequent answering: $a=\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle).$
@@ 5 Experiments
Method NQ TriviaQA BioASQ PopQA Overall NQ TriviaQA BioASQ PopQA Overall
Claude 3.5 Sonnet (20240620) Gemini 1.5 Pro (002)
No RAG 47.1 82.0 50.4 29.8 54.5 44.8 80.2 45.8 25.3 51.3
RAG 44.4 76.7 58.0 36.0 55.5 42.7 76.0 55.2 33.7 53.7
USC (Chen et al., 2024b) 48.1 80.2 61.5 37.6 58.7 46.4 76.7 58.4 37.6 56.4
GenRead (Yu et al., 2023a) 42.0 74.2 57.0 34.3 53.6 45.1 77.4 54.9 34.3 54.7
RobustRAG (Xiang et al., 2024) 47.8 78.1 56.3 37.1 56.5 34.2 67.5 44.1 32.0 45.6444We observe a high refusal rate in RobustRAG for Gemini.
InstructRAG (Wei et al., 2024) 47.1 83.0 58.0 41.0 58.8 46.8 80.6 54.9 34.8 56.1
Self-Route (Xu et al., 2024) 47.5 78.8 59.1 41.0 58.1 47.5 79.9 58.0 38.2 57.6
Astute RAG 52.2 84.1 60.1 44.4 61.7 50.2 81.6 58.0 40.5 59.2
Mistral-Large (2407), 128B Mistral-Nemo (2407), 12B
No RAG 46.8 79.5 43.7 24.7 51.1 29.8 67.8 34.3 23.0 40.2
RAG 43.1 77.4 55.9 36.0 54.7 39.3 66.8 49.0 32.6 48.3
USC (Chen et al., 2024b) 51.2 80.9 61.5 36.0 59.5 29.5 66.1 36.0 20.2 39.6
GenRead (Yu et al., 2023a) 40.7 73.1 55.6 35.4 52.7 38.6 68.9 48.3 33.7 48.7
RobustRAG (Xiang et al., 2024) 42.7 77.7 50.4 34.8 53.0 35.6 71.7 44.1 27.5 46.4
InstructRAG (Wei et al., 2024) 45.4 80.6 57.3 36.5 56.7 38.3 61.8 50.4 23.6 45.5
Self-Route (Xu et al., 2024) 45.4 77.7 57.3 38.2 56.2 41.4 73.5 51.8 30.9 51.2
Astute RAG 50.2 82.7 58.4 42.1 59.9 42.7 73.9 49.3 32.6 51.3
Table 1: Main results on Claude 3.5 Sonnet, Gemini 1.5 Pro, Mistral-Large, and Mistral-Nemo under zero-shot setting, showing the accuracy of benchmarked alternatives vs. Astute RAG. Best scores are in bold. Note that USC consumes approximately three times more tokens than other RAG methods, and is not directly comparable.
We evaluate the effectiveness of Astute RAG on overcoming imperfect retrieval augmentation and addressing knowledge conflicts. In this section, we introduce the experiment setting (Sec. 5.1), compare the performance of Astute RAG with various baselines on diverse datasets (Sec. 5.2), and provide in-depth analyses (Sec. 5.3).
@@ 5.1 Experimental Settings
Datasets and metrics.
We consider datasets encompass general questions, domain-specific questions, long-tail questions, as well as both short-form and long-form formats, following prior work Xiang et al. (2024); Wei et al. (2024). On NQ, TriviaQA, BioASQ, and PopQA, we provide 10 passages collected with Google Search from the Web for each instance. For long-form QA, we use ASQA Stelmakh et al. (2022). We also evaluate on RGB (Chen et al., 2024a). We choose the English subset (refined version) focusing on noise robustness. For each instance, we select five top negative passages to form a worst-case scenario. Following prior work, we report the accuracy by string match. More details are in Appx. B.
Models and General Settings.
@@
We conduct experiments on advanced proprietary and open-source LLMs of different scales, including Claude 3.5 Sonnet (claude-3-5-sonnet@20240620), [^5] https://www.anthropic.com/news/claude-3-5-sonnet Gemini 1.5 Pro (gemini-1.5-pro-002), [^6] https://deepmind.google/technologies/gemini/pro/ Mistral-Large (128B; version 2407), and Mistral-Nemo (12B; version 2407). The generation temperature is set to 0 and the maximum output tokens is set to 1,024. All experiments are under the zero-shot setting for controlled evaluation.
Baselines.
@@
We compare Astute RAG with various RAG methods designed for enhanced robustness. USC (Chen et al., 2024b) is a self-consistency method that samples multiple LLM responses and aggregates the answers. It provides a reference of naive improvements using additional API calls. Genread (Yu et al., 2023a) augments retrieved passages with LLM-generated passages without explicit consolidation process. RobustRAG (Xiang et al., 2024) aggregates answers from independent passages to provide certifiable robustness. We use the best-performing keyword aggregation variant. InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages. For a fair comparison, no training is applied. Self-Route (Xu et al., 2024) adaptively switches between LLMs with and without RAG. [^7] The original Self-Route switches between RAG and long-context LLMs, while our implementation switches between RAG and No RAG according to our problem formulation. It provides a reference of switching between LLMs’ internal and external knowledge.
Implementation Details.
The prompt templates for Astute RAG can be found in Appx. A. By default, we set $t=1$ and $\hat{m}=1$ to limit the number of additional tokens used. Results with larger $t$ and $\hat{m}$ are discussed in Sec. 5.3.
Figure 5: Performance on ASQA.
Figure 6: Worst-case performance of Claude on RGB. Astute RAG reaches a performance close to No RAG, while other RAG systems are far behind.
@@ 5.2 Main Results
Performance under real-world retrieval.
Tab. 1 presents the results with real-world retrieval augmentation of various LLMs. We find that retrieved passages might not always bring benefits – on NQ and TriviaQA, RAG performance lags behind No RAG for advanced LLMs. We attribute this questions being covered by the LLM’s internal knowledge and the noise in retrieval results misleading the LLM. In contrast, on BioASQ and PopQA, which focus on domain-specific and ‘long-tail’ questions, RAG significantly improves the LLM performance. Due to imperfect retrieval augmentation, however, the absolute performance still remains to be unsatisfactory. Among all baselines, no single method consistently outperforms others across all datasets and LLMs. This observation highlights these baselines being tailored to distinct settings and not being universally applicable. Overall, InstructRAG and Self-Route demonstrate relatively superior performance among other alternatives. Astute RAG consistently outperforms baselines across all LLMs in terms of overall accuracy. The relative improvement compared to the best baseline is 6.85% for Claude and 4.13% for Gemini, with the improvements in domain-specific questions being much higher. These highlight the effectiveness of Astute RAG in overcoming imperfect retrieval augmentation and knowledge conflicts. Additionally, we observe consistent improvements on the open-source Mistral models. The results demonstrate that Astute RAG generalizes well to LLMs of smaller sizes.
Figure 7: Performance across different retrieval precision buckets. Astute RAG is consistently better.
Performance on long-form QA.
We conduct additional experiments on the long-form QA dataset, ASQA. Fig. 5 demonstrates that Astute RAG consistently achieves significant improvements, reinforcing its effectiveness across diverse scenarios.
Worst-case performance on RGB.
Fig. 6 presents the results under the worst-case setting on RGB where all retrieved documents are negative, to demonstrate robustness. The performance gap between RAG and No RAG exceeds 50 points, highlighting the detrimental impact of imperfect retrieval results and emphasizing the importance of providing robust safeguards against worst-case scenarios. While the baseline RAG methods outperform the original RAG, they still obviously fall behind ‘No RAG’. Astute RAG is the only RAG method that reaches a performance close to ‘No RAG’, further supporting its effectiveness in addressing imperfect retrieval augmentation.
@@ 5.3 Analyses
We conduct in-depth analyses using Claude following the setting of Tab. 1.
The impact of retrieval precision.
@@
As shown in Fig. 7, Astute RAG achieves consistently better performance across different retrieval precision regimes, indicating its effectiveness in improving RAG trustworthiness in broad scenarios. Notably, Astute RAG does not sacrifice performance gain under high retrieval quality in exchange for improvement under low retrieval quality. When the retrieval quality is extremely low (close to zero precision), all other RAG variants underperform the ’No RAG’ baseline, except for Astute RAG.
Figure 8: Performance on conflicting and consistent instances between No RAG and RAG.
Addressing knowledge conflicts.
@@
We split our collected data into three subset according to the answers with and without RAG: the answers from two can be (i) both correct, (ii) both incorrect, or (iii) conflicting with one being correct. The results are shown in Fig. 8. On the conflicting subset, Astute RAG successfully chooses the correct answer in approximately 80% of cases, being the most effective one in addressing knowledge conflicts. Notably, Astute RAG even brings performance improvement on the subset where neither internal nor external knowledge alone leads to the correct answer. This indicates that Astute RAG can effectively combine partially-correct information from LLM-internal and external knowledge.
Benefits of more consolidation iteration.
For efficiency, we employ a single iteration of knowledge consolidation in our main experiments. However, incorporating multiple iterations has the potential to further enhance model performance as shown in Fig. 9. The magnitude of this improvement diminishes as $t$ increases, indicating that the knowledge has been better presented and less improvement space left after each iteration.
Efficiency in tokens consumed and API calls.
As a proxy to overall prediction cost and latency, we present the average number of tokens and API calls used per instance in Fig. 10 and Fig. 11. Astute RAG incurs only a marginal cost increase, <5%, while delivering substantial improvement, >11%, compared to the RAG baseline.
Figure 9: Accuracy improvement when increasing $t$.
Effectiveness of adaptive generation.
The results in Tab. 2 illustrate the model’s performance when varying the maximum number of passages generated. The design of adaptive generation has been effectively reflected, as the number of generated passages is dynamically adjusted leading to $m<\hat{m}$. Notably, the number of generated passages can be controlled by $\hat{m}$, and results show that the system does not generate passages excessively.
Impact of Source-Awareness.
To evaluate the impact of source-awareness, we conducted an ablation study where source labels (own memory and external retrieval) were removed during the consolidation and answer generation process. The results are shown in Tab. 3. The comparison shows that providing source information (Astute RAG) leads to better performance overall compared to omitting it (Astute RAG (No Source)), particularly on NQ, PopQA, and BioASQ, suggesting that awareness of information origin aids the consolidation process.
Accuracy of intermediate steps.
To investigate the performance of intermediate steps, including knowledge consolidation and confidence assignment, we use LLM-as-a-judge with the instruction in Appx. A. Our experimental results show that the accuracy for knowledge consolidation is 98.2%, and for confidence assignment, it is 95.0%. These results demonstrate the effectiveness of the proposed framework in the intermediate stages.
Qualitative examples.
@@
In Fig. 12, we present two representative examples showing the intermediate outputs of Astute RAG. In the first example, LLM without RAG generates a wrong answer, while RAG returns a correct answer. Astute RAG successfully identified the incorrect information in its generated passage and an external passage, avoiding confirmation bias Tan et al. (2024). In the second example, LLM is correct but RAG is incorrect due to imperfect retrieval. Astute RAG detected the correct answer from imperfect context leveraging internal knowledge.
Figure 10: Efficiency in terms of tokens consumed.